OpenStudy (anonymous):

Find the velocity, v(t), for an object moving along the x-axis if the acceleration, a(t), is a(t) = 2t + sin(t) and v(0) = 4. a: v(t) = t^2 + cos(t) + 3 b: v(t) = 2 + cos(t) + 1 c: v(t) = t^2 − cos(t) + 5 d: v(t) = t^2 + sin(t) + 4

OpenStudy (solomonzelman):

take the integral of 2t+sin(t)

OpenStudy (solomonzelman):

for your "C" value use the fact that v(0)=4, to determine what the "c" will be

OpenStudy (anonymous):

okay the integral is: t^2-cos(t)

OpenStudy (solomonzelman):

`\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}2t+\sin(t)~dt}\)` you can use my latex.

OpenStudy (solomonzelman):

yes,

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}2t+\sin(t)~dt~=~t^2-\cos(t)+C}\)

OpenStudy (anonymous):

I think i got it! is it B? :)

OpenStudy (anonymous):

V(t)=2+cos(t)+1 V(0)=2+1+1 V(0)=4

OpenStudy (solomonzelman):

and you know that \(\large\color{slate}{\displaystyle~v(\color{red}{0})=\color{blue}{4}}\), saying that \(\large\color{slate}{\displaystyle~(\color{red}{0})^2-\cos(\color{red}{0})+C=\color{blue}{4}}\)

OpenStudy (solomonzelman):

the velocity function is an integral of acceleration

OpenStudy (anonymous):

Ohh then c = 5

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

and your answer is?

OpenStudy (solomonzelman):

you can already deduce by the fact the we have the cos(y) since that is integral of sin(t), and that it is negative cos(t).

OpenStudy (anonymous):

:) c: v(t) = t^2 − cos(t) + 5 thank u!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (solomonzelman):

I mean cos(t)

OpenStudy (solomonzelman):

yes C is correct

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle~v(t)=t^2-\cos(t)+5}\)

OpenStudy (solomonzelman):

you welcome !

OpenStudy (anonymous):

Ur so good!!!!!!!!!!!!!

OpenStudy (solomonzelman):

that is debatable, but good that I had that impression.

OpenStudy (solomonzelman):

yw, and I will definitely take the complement. tnx

OpenStudy (anonymous):

I have another question, can u plz help me with it? :)